# Transfer of quantum game strategies

### v1

@@ 4. SQNS correlations
(12) ${\rm Tr}_{X_{1}}\Gamma(\rho_{X_{2}Y_{2}}\otimes\rho_{A_{1}B_{1}})=0\text{ if }\rho_{X_{2}Y_{2}}\in M_{X_{2}Y_{2}}\text{ and }{\rm Tr}_{X_{2}}(\rho_{X_{2}Y_{2}})=0,$
(13) ${\rm Tr}_{Y_{1}}\Gamma(\rho_{X_{2}Y_{2}}\otimes\rho_{A_{1}B_{1}})=0\text{ if }\rho_{X_{2}Y_{2}}\in M_{X_{2}Y_{2}}\text{ and }{\rm Tr}_{Y_{2}}(\rho_{X_{2}Y_{2}})=0,$
(14) ${\rm Tr}_{A_{2}}\Gamma(\rho_{X_{2}Y_{2}}\otimes\rho_{A_{1}B_{1}})=0\text{ if }\rho_{A_{1}B_{1}}\in M_{A_{1}B_{1}}\text{ and }{\rm Tr}_{A_{1}}(\rho_{A_{1}B_{1}})=0,$
@@
(15) ${\rm Tr}_{B_{2}}\Gamma(\rho_{X_{2}Y_{2}}\otimes\rho_{A_{1}B_{1}})=0\text{ if }\rho_{A_{1}B_{1}}\in M_{A_{1}B_{1}}\text{ and }{\rm Tr}_{B_{1}}(\rho_{A_{1}B_{1}})=0.$
@@
$\sum_{x_{1}\in X_{1}}p(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}b_{1})=\sum_{x_{1}\in X_{1}}p(x_{1}y_{1},a_{2}b_{2}|x_{2}^{\prime}y_{2},a_{1}b_{1}),\ \ x_{2},x_{2}^{\prime}\in X_{2},$
$\sum_{y_{1}\in Y_{1}}p(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}b_{1})=\sum_{y_{1}\in Y_{1}}p(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2}^{\prime},a_{1}b_{1}),\ \ y_{2},y_{2}^{\prime}\in Y_{2},$
$\sum_{a_{2}\in A_{2}}p(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}b_{1})=\sum_{a_{2}\in A_{2}}p(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}^{\prime}b_{1}),\ \ a_{1},a_{1}^{\prime}\in A_{1},$
@@
$\sum_{b_{2}\in B_{2}}p(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}b_{1})=\sum_{b_{2}\in B_{2}}p(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}b_{1}^{\prime}),\ \ b_{1},b_{1}^{\prime}\in B_{1}.$
@@
(16) $\sum\limits_{x_{2}\in X_{2}}\sum\limits_{y_{2},y_{2}^{\prime}\in Y_{2}}\tilde{\rho}_{x_{2}.x_{2}}^{y_{2},y_{2}^{\prime}}\epsilon_{y_{2}y_{2}^{\prime}}=0.$
@@
×$\;\;\;\;\;\;\;\;\times\epsilon_{y_{1}y_{1}}\otimes\epsilon_{a_{2}a_{2}}\otimes\epsilon_{b_{2}b_{2}}$
@@
×$\;\;\;\;\;\;\;\;\times\epsilon_{y_{1}y_{1}}\otimes\epsilon_{a_{2}a_{2}}\otimes\epsilon_{b_{2}b_{2}}$
$=$ $0$
@@
×$\;\;\;\;\;\;\;\;\;\;\;\;\times\epsilon_{x_{1}x_{1}}\otimes\epsilon_{y_{1}y_{1}}\otimes\epsilon_{a_{2}a_{2}}\otimes\epsilon_{b_{2}b_{2}}.$
@@
(17) $\Phi^{*}(1\otimes M_{Y_{1}}\otimes M_{A_{2}}\otimes M_{B_{2}})\subseteq 1\otimes M_{Y_{2}}\otimes M_{A_{1}}\otimes M_{B_{1}},$
(18) $\Phi^{*}(M_{X_{1}}\otimes 1\otimes M_{A_{2}}\otimes M_{B_{2}})\subseteq M_{X_{2}}\otimes 1\otimes M_{A_{1}}\otimes M_{B_{1}},$
(19) $\Phi^{*}(M_{X_{1}}\otimes M_{Y_{1}}\otimes 1\otimes M_{B_{2}})\subseteq M_{X_{2}}\otimes M_{Y_{2}}\otimes 1\otimes M_{B_{1}},$
@@
(20) $\Phi^{*}(M_{X_{1}}\otimes M_{Y_{1}}\otimes M_{A_{2}}\otimes 1)\subseteq M_{X_{2}}\otimes M_{Y_{2}}\otimes M_{A_{1}}\otimes 1.$
@@
We denote by $\mathcal{Q}_{\rm sqc}$ (resp. $\mathcal{Q}_{\rm sqa},\mathcal{Q}_{\rm sq},\mathcal{Q}_{\rm sloc}$) the classes of quantum commuting (resp. approximately quantum, quantum, local) SQNS correlations. We note that, by definition, $\mathcal{Q}_{\rm st}\subseteq\mathcal{Q}_{\rm t}$, for ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$. In the sequel, for ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$ let $\mathcal{C}_{\rm st}$ denote the class of all SNS correlations of type ${\rm t}$, as defined in [19, Section 5].
@@
We pause here for a correction to a previous work, and some necessary clarifying remarks. In [19], we defined a classical correlation $\Gamma$ to belong to the class $\mathcal{C}_{\rm sqc}$ if there existed a Hilbert space $H$, dilatable NS operator matrices $P=(P_{x_{2}y_{2},x_{1}y_{1}})_{x_{2}y_{2},x_{1}y_{1}}$ and $Q=(Q_{a_{1}b_{1},a_{2}b_{2}})_{a_{1}b_{1},a_{2}b_{2}}$ on $H$ with mutually commuting entries, and a unit vector $\xi\in H$ such that $\Gamma=\Gamma_{P,Q,\xi}$. In [19, Lemma 5.8] we claimed that this was a necessary and sufficient condition for the existence of a Hilbert space $K$, PVM’s $(P_{x_{2},x_{1}})_{x_{1}\in X_{1}},(P^{y_{2},y_{1}})_{y_{1}\in Y_{1}},(Q_{a_{1},a_{2}})_{a_{2}\in A_{2}}$ and $(Q^{b_{1},b_{2}})_{b_{2}\in B_{2}}$ on $K$ with mutually commuting entries, and a unit vector $\eta\in K$ such that
@@
for all $x_{i}\in X_{i},y_{i}\in Y_{i},a_{i}\in A_{i},b_{i}\in B_{i},i=1,2$. However, we have since discovered a gap in the proof of said lemma resulting from the incorrect use of associativity of the commuting operator system tensor product.
In generalizing to the case of what we call SQNS correlations, as the natural generalization of NS operator matrices are strongly stochastic operator matrices if we were to follow the path laid out in [19, Definition 5.6 (i)] one would expect that we say $\Gamma\in\mathcal{Q}_{\rm sqc}$ if there exists a Hilbert space $H$, dilatable strongly stochastic operator matrices $P=(P_{x_{2}x_{2}^{\prime},y_{2}y_{2}^{\prime}}^{x_{1}x_{1}^{\prime},y_{1}y_{1}^{\prime}})_{x_{2}x_{2}^{\prime},y_{2}y_{2}^{\prime},x_{1}x_{1}^{\prime},y_{1}y_{1}^{\prime}}$ and $Q=(Q_{a_{1}a_{1}^{\prime},b_{1}b_{1}^{\prime}}^{a_{2}a_{2}^{\prime},b_{2}b_{2}^{\prime}})_{a_{1}a_{1}^{\prime},b_{1}b_{1}^{\prime},a_{2}a_{2}^{\prime},b_{2}b_{2}^{\prime}}$ with mutually commuting entries acting on $H$, and a unit vector $\xi\in H$ such that $\Gamma=\Gamma_{P,Q,\xi}$. However, in order to get the decomposition of strongly stochastic operator matrices that we will need for the results in Section 5 due to the failure of [19, Lemma 5.8] (and its extension) we must instead require it in the definition (as seen in Definition 4.4 (i)). Furthermore, we will need to amend the definition for $\Gamma\in\mathcal{C}_{\rm sqc}$, which we do below. We point out that working with Definition 4.6 (see below) instead of [19, Definition 5.6(i)] recovers all subsequent results in [19] which were instead obtained using [19, Lemma 5.8].
@@
In the case when ${\rm t}={\rm ns}$, this follows directly by Remark 4.1. For ${\rm t}={\rm loc,q,qa}$ or ${\rm qc}$, this follows from Remark 3.2 and [38, Lemma 7.2]. ∎
@@
(21) $\Gamma=\sum\limits_{j=1}^{k}\lambda_{j}\Phi_{X}^{(j)}\otimes\Phi_{Y}^{(j)}\otimes\Phi_{A}^{(j)}\otimes\Phi_{B}^{(j)},$

### v2

@@ 3.1. SQNS correlations
${\rm Tr}_{X_{1}}\Gamma(\rho_{X_{2}Y_{2}}\otimes\rho_{A_{1}B_{1}})=0\text{ if }\rho_{X_{2}Y_{2}}\in M_{X_{2}Y_{2}}\text{ and }{\rm Tr}_{X_{2}}(\rho_{X_{2}Y_{2}})=0,$ (12)
${\rm Tr}_{Y_{1}}\Gamma(\rho_{X_{2}Y_{2}}\otimes\rho_{A_{1}B_{1}})=0\text{ if }\rho_{X_{2}Y_{2}}\in M_{X_{2}Y_{2}}\text{ and }{\rm Tr}_{Y_{2}}(\rho_{X_{2}Y_{2}})=0,$ (13)
${\rm Tr}_{A_{2}}\Gamma(\rho_{X_{2}Y_{2}}\otimes\rho_{A_{1}B_{1}})=0\text{ if }\rho_{A_{1}B_{1}}\in M_{A_{1}B_{1}}\text{ and }{\rm Tr}_{A_{1}}(\rho_{A_{1}B_{1}})=0,$ (14)
@@
${\rm Tr}_{B_{2}}\Gamma(\rho_{X_{2}Y_{2}}\otimes\rho_{A_{1}B_{1}})=0\text{ if }\rho_{A_{1}B_{1}}\in M_{A_{1}B_{1}}\text{ and }{\rm Tr}_{B_{1}}(\rho_{A_{1}B_{1}})=0.$ (15)
@@ A $\it classical\;strongly\;no-signalling\;(SNS)\;correlation$} [20] is a correlation
$\sum_{x_{1}\in X_{1}}\hskip-2.84544ptp(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}b_{1})\hskip-1.42271pt=\hskip-5.69046pt\sum_{x_{1}\in X_{1}}\hskip-2.84544ptp(x_{1}y_{1},a_{2}b_{2}|x_{2}^{\prime}y_{2},a_{1}b_{1}),\ \ x_{2},x_{2}^{\prime}\in X_{2},$
$\sum_{y_{1}\in Y_{1}}\hskip-2.84544ptp(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}b_{1})\hskip-1.42271pt=\hskip-5.69046pt\sum_{y_{1}\in Y_{1}}\hskip-2.84544ptp(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2}^{\prime},a_{1}b_{1}),\ \ y_{2},y_{2}^{\prime}\in Y_{2},$
$\sum_{a_{2}\in A_{2}}\hskip-2.84544ptp(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}b_{1})\hskip-1.42271pt=\hskip-5.69046pt\sum_{a_{2}\in A_{2}}\hskip-2.84544ptp(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}^{\prime}b_{1}),\ \ a_{1},a_{1}^{\prime}\in A_{1},$
@@
$\sum_{b_{2}\in B_{2}}\hskip-2.84544ptp(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}b_{1})\hskip-1.42271pt=\hskip-5.69046pt\sum_{b_{2}\in B_{2}}\hskip-2.84544ptp(x_{1}y_{1},a_{2}b_{2}|x_{2}y_{2},a_{1}b_{1}^{\prime}),\ \ b_{1},b_{1}^{\prime}\in B_{1}.$
@@
$\sum\limits_{x_{2}\in X_{2}}\sum\limits_{y_{2},y_{2}^{\prime}\in Y_{2}}\tilde{\rho}_{x_{2}.x_{2}}^{y_{2},y_{2}^{\prime}}\epsilon_{y_{2}y_{2}^{\prime}}=0.$ (16)
@@
$\;\;\;\;\;\;\;\;\times\epsilon_{y_{1}y_{1}}\otimes\epsilon_{a_{2}a_{2}}\otimes\epsilon_{b_{2}b_{2}}$
@@
$\;\;\;\;\;\;\;\;\times\epsilon_{y_{1}y_{1}}\otimes\epsilon_{a_{2}a_{2}}\otimes\epsilon_{b_{2}b_{2}}$
$=$ $0.$
@@
$\;\;\;\;\;\;\;\;\;\;\;\;\times\epsilon_{x_{1}x_{1}}\otimes\epsilon_{y_{1}y_{1}}\otimes\epsilon_{a_{2}a_{2}}\otimes\epsilon_{b_{2}b_{2}}.$
@@
$\Phi^{*}(1\otimes M_{Y_{1}}\otimes M_{A_{2}}\otimes M_{B_{2}})\subseteq 1\otimes M_{Y_{2}}\otimes M_{A_{1}}\otimes M_{B_{1}},$ (17)
$\Phi^{*}(M_{X_{1}}\otimes 1\otimes M_{A_{2}}\otimes M_{B_{2}})\subseteq M_{X_{2}}\otimes 1\otimes M_{A_{1}}\otimes M_{B_{1}},$ (18)
$\Phi^{*}(M_{X_{1}}\otimes M_{Y_{1}}\otimes 1\otimes M_{B_{2}})\subseteq M_{X_{2}}\otimes M_{Y_{2}}\otimes 1\otimes M_{B_{1}},$ (19)
@@
$\Phi^{*}(M_{X_{1}}\otimes M_{Y_{1}}\otimes M_{A_{2}}\otimes 1)\subseteq M_{X_{2}}\otimes M_{Y_{2}}\otimes M_{A_{1}}\otimes 1.$ (20)
@@
We denote by $\mathcal{Q}_{\rm sqc}$ (resp. $\mathcal{Q}_{\rm sqa},\mathcal{Q}_{\rm sq},\mathcal{Q}_{\rm sloc}$) the classes of quantum commuting (resp. approximately quantum, quantum, local) SQNS correlations. We note that, by definition, $\mathcal{Q}_{\rm st}\subseteq\mathcal{Q}_{\rm t}$, for ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$. In the sequel, for ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$ let $\mathcal{C}_{\rm st}$ denote the class of all SNS correlations of type $\rm t$}, as defined in [20, Section 5].
@@
We pause here for a correction to a previous work, and some necessary clarifying remarks. In [20], we defined a classical correlation $\Gamma$ to belong to the class $\mathcal{C}_{\rm sqc}$ if there existed a Hilbert space $H$, dilatable NS operator matrices $P=(P_{x_{2}y_{2},x_{1}y_{1}})_{x_{2}y_{2},x_{1}y_{1}}$ and $Q=(Q_{a_{1}b_{1},a_{2}b_{2}})_{a_{1}b_{1},a_{2}b_{2}}$ on $H$ with mutually commuting entries, and a unit vector $\xi\in H$ such that $\Gamma=\Gamma_{P,Q,\xi}$. In [20, Lemma 5.8] we claimed that this was a necessary and sufficient condition for the existence of a Hilbert space $K$, PVM’s $(P_{x_{2},x_{1}})_{x_{1}\in X_{1}},(P^{y_{2},y_{1}})_{y_{1}\in Y_{1}},(Q_{a_{1},a_{2}})_{a_{2}\in A_{2}}$ and $(Q^{b_{1},b_{2}})_{b_{2}\in B_{2}}$ on $K$ with mutually commuting entries, and a unit vector $\eta\in K$ such that
@@
for all $x_{i}\in X_{i},y_{i}\in Y_{i},a_{i}\in A_{i},b_{i}\in B_{i},i=1,2$. However, we have since discovered a gap in the proof of said lemma resulting from the incorrect use of associativity of the commuting operator system tensor product (see Section 6).
In generalizing to the case of what we call SQNS correlations, as the natural generalization of NS operator matrices are strongly stochastic operator matrices if we were to follow the path laid out in [20, Definition 5.6 (i)] one would expect that we say $\Gamma\in\mathcal{Q}_{\rm sqc}$ if there exists a Hilbert space $H$, dilatable strongly stochastic operator matrices $P=(P_{x_{2}x_{2}^{\prime},y_{2}y_{2}^{\prime}}^{x_{1}x_{1}^{\prime},y_{1}y_{1}^{\prime}})_{x_{2}x_{2}^{\prime},y_{2}y_{2}^{\prime},x_{1}x_{1}^{\prime},y_{1}y_{1}^{\prime}}$ and $Q=(Q_{a_{1}a_{1}^{\prime},b_{1}b_{1}^{\prime}}^{a_{2}a_{2}^{\prime},b_{2}b_{2}^{\prime}})_{a_{1}a_{1}^{\prime},b_{1}b_{1}^{\prime},a_{2}a_{2}^{\prime},b_{2}b_{2}^{\prime}}$ with mutually commuting entries acting on $H$, and a unit vector $\xi\in H$ such that $\Gamma=\Gamma_{P,Q,\xi}$. However, in order to get the decomposition of strongly stochastic operator matrices that we will need for the results in Section 6 due to the failure of [20, Lemma 5.8] (and its extension) we must instead require it in the definition (as seen in Definition 3.7 (i)). Furthermore, we will need to amend the definition for $\Gamma\in\mathcal{C}_{\rm sqc}$, which we do below. We point out that working with Definition 3.9 (see below) instead of [20, Definition 5.6(i)] recovers all subsequent results in [20] which were instead obtained using [20, Lemma 5.8].
@@
In the case when ${\rm t}={\rm ns}$, this follows directly by Remark 3.4. For ${\rm t}={\rm loc,q,qa}$ or $\rm qc$}, this follows from Remark 3.2 and [39, Lemma 7.2]. ∎
@@
$\Gamma=\sum\limits_{j=1}^{k}\lambda_{j}\Phi_{X}^{(j)}\otimes\Phi_{Y}^{(j)}\otimes\Phi_{A}^{(j)}\otimes\Phi_{B}^{(j)},$ (21)